OpenStudy (anonymous):

Solve the linear system. 5x+6y=-9 -x-8y=-5

OpenStudy (anonymous):

multiply the second equation by \(5\) then add them up in a column

OpenStudy (anonymous):

On both sides?

OpenStudy (anonymous):

yes of course multiply every part of the second equation by 5 the \(-x\) the \(-8y\) and also the \(-5\)

OpenStudy (anonymous):

Okay I got y=1. Is that it?

OpenStudy (anonymous):

idk i didn't do it

OpenStudy (anonymous):

Well do you not need to solve for x in a linear system?

OpenStudy (anonymous):

yes it is \(y=1\) right

OpenStudy (anonymous):

you know \(y=1\) so you certainly know what \(x\) is because \( 5x+6y=-9\) and since \(y=1\) this says \[5x+6\times 1=-9\\ 5x+6=-9\] etc

OpenStudy (anonymous):

Oh I see. Thanks . . once again. xD